# Efficient Learning with Sine-Activated Low-Rank Matrices

### v4

@@ 6 Conclusion
In this work we have demonstrated that applying a sinusoidal non-linearity improves the accuracy of low-rank approximations by increasing their rank. While simple, this method is highly applicable to parameter constrained models such as LoRA, as it improves approximation without adding capacity, by overcoming representation limits of the matrix rank. We have fully justified this approach from theoretical first principles. When applied as a drop-in component we showed that this method leads to surprisingly large improvements across a range of tasks involving low-rank models, including language tasks, image classification, neural radiance fields and 3D shape modelling.
@@ A.1 Theoretical Framework
Observer by definition of the Frobenius norm that
@@
Observe that from Eqn. equation 25 it suffices to prove the lower bound on $SR(\mathbf{A})$. This is immediate from lemma 1 and lemma 2. ∎
@@ A.2 Experiments
For the experiments we observed that the gain factor $g$ in Equation 3 should be chosen analogously to how weights are initialized in (He et al., 2015a). In particular we chose $g=\sqrt{n}$, where $n$ was the number of rows of the weight matrix. During backpropagation the frequency parameter $\omega$ scales the gradients which can cause gradient exploding. To mitigate this we found the choice of $g=\sqrt{n}$ worked best. We also point out that for the experiments there is no principled way to set $\omega$. Therefore, we will obtain $\omega$ by treating it as a hyperparameter and tuning it according to what gives the best results.
@@ A.2.1 Fine tuning large language models
We followed the settings in (Liu et al., 2024). In the Transformer architecture, there are four weight matrices in the self-attention module ($\mathbf{W_{\text{q}}},\mathbf{W_{\text{k}}},\mathbf{W_{\text{v}}},\mathbf{W_{\text{o}}}$) and two in the MLP module($\mathbf{W_{\text{up}}},\mathbf{W_{\text{down}}}$). To evaluate Llama3-8B, we implement low-rank adaptation only on $\mathbf{W_{\text{q}}},\mathbf{W_{\text{k}}},\mathbf{W_{\text{v}}},\mathbf{W_{\text{up}}},\;\text{and}\;\mathbf{W_{\text{down}}}$. We study the performance of LoRA and sine LoRA in terms of different rank $k={4,8,16,32}$ and configurations are as shown in 5.

### v5

@@ 6 Conclusion
In this work we have demonstrated that applying a sinusoidal non-linearity improves the accuracy of low-rank approximations by increasing their rank. While simple, this method is highly applicable to parameter-constrained models such as LoRA, as it improves approximation without adding capacity, by overcoming representation limits of the matrix rank. We have fully justified this approach from theoretical first principles. When applied as a drop-in component we showed that this method leads to surprisingly large improvements across a range of tasks involving low-rank models, including language tasks, image classification, neural radiance fields and 3D shape modelling.
Acknowledgements
Hemanth Saratchandran and Simon Lucey acknowledge support from Commonwealth Bank of Australia through the CommBank Centre for Foundational AI Research. This funding was essential for the completion of the research described in this publication.
@@ A.1 Theoretical Framework
Observe by definition of the Frobenius norm that
@@
Observe that from equation 25 it suffices to prove the lower bound on $SR(\mathbf{A})$. This is immediate from lemma 1 and lemma 2. ∎
@@ A.2 Experiments
For the experiments we observed that the gain factor $g$ in Equation 3 should be chosen analogously to how weights are initialized in (He et al., 2015). In particular we chose $g=\sqrt{n}$, where $n$ was the number of rows of the weight matrix. During backpropagation the frequency parameter $\omega$ scales the gradients which can cause gradient exploding. To mitigate this we found the choice of $g=\sqrt{n}$ worked best. We also point out that for the experiments there is no principled way to set $\omega$. Therefore, we will obtain $\omega$ by treating it as a hyperparameter and tuning it according to best results.
@@ A.2.1 Fine tuning large language models
We followed the settings in (Liu et al., 2024). In the Transformer architecture, there are four weight matrices in the self-attention module ($\mathbf{W_{\text{q}}},\mathbf{W_{\text{k}}},\mathbf{W_{\text{v}}},\mathbf{W_{\text{o}}}$) and two in the MLP module($\mathbf{W_{\text{up}}},\mathbf{W_{\text{down}}}$). To evaluate Llama3-8B, we implement low-rank adaptation only on $\mathbf{W_{\text{q}}},\mathbf{W_{\text{k}}},\mathbf{W_{\text{v}}},\mathbf{W_{\text{up}}},\;\text{and}\;\mathbf{W_{\text{down}}}$. We study the performance of LoRA and sine LoRA for different rank $k={4,8,16,32}$ and configurations are as shown in 5.